### v1

# Duet: Agentic Design Understanding via Experimentation and Testing

Abstract
AI agents powered by large language models (LLMs) are being used to solve increasingly complex software engineering challenges, but struggle with hardware design tasks. Register Transfer Level (RTL) code presents a unique challenge for LLMs, as it encodes complex, dynamic, time-evolving behaviors using the low-level language features of SystemVerilog. LLMs struggle to infer these complex behaviors from the syntax of RTL alone, which limits their ability to complete all downstream tasks like code completion, documentation, or verification. In response to this issue, we present Duet: a general methodology for developing Design Understanding via Experimentation and Testing. Duet mimics how hardware design experts develop an understanding of complex designs: not just via a one-off readthrough of the RTL, but via iterative experimentation using a number of tools. Duet iteratively generates hypotheses, tests them with EDA tools (e.g., simulation, waveform inspection, and formal verification), and integrates the results to build a bottom-up understanding of the design. In our evaluations, we show that Duet improves AI agent performance on formal verification, when compared to a baseline flow without experimentation.
I Introduction
AI agents are quickly taking over many tasks in software engineering and beyond. Powered by large language models (LLMs), AI agents can be deployed as powerful autonomous software workflows. An AI agent takes a text prompt as input, and is generally also equipped with tools it can call (e.g., Python functions or command-line utilities to read and write files, execute computations, or run installed tools). The agent then iteratively queries the LLM (starting with the prompt) to get the next action. When the LLM requests an action, the function or command-line utility is called and the results are sent back to the LLM. Eventually, the LLM decides to stop (or is stopped externally), and sends a final result. Throughout the process, the agent may have also created a number of artifacts, such as files in the filesystem. Using this general structure, AI agents have been able to replicate many human tasks.
While AI agents have shown human-level performance in software engineering tasks, they continue to struggle with similar tasks in hardware design. Among many reasons, we hypothesize that this is because the underlying LLMs powering AI-assisted tools struggle with Register Transfer Level (RTL) code. RTL is inherently complex, often capturing dynamic, time-evolving behaviors using the low-level, heavily implicit syntax of languages like SystemVerilog. This is in stark contrast to software languages like Python or Java which contain more structure in the syntax of the code itself; for example, sequential lines in imperative programming languages generally correspond to instructions which will execute over time, and function names can be used to understand when control flow jumps across files in the codebase. On the other hand, RTL does not have such structure. For example, sequential states in a finite state machine (FSM) might be separated by tens or hundreds of lines in a case statement, and the control flow between these states may be much less explicit. Thus, when LLMs are given only the RTL, they struggle to understand the deeper behaviors of hardware designs and build design understanding. As a result, AI agents perform worse on downstream hardware tasks like verification or debugging.
However, hardware designers themselves do not build their understanding of a design simply by reading the RTL. Instead, the process is more dynamic; designers will use tools like simulation, waveform viewers, and formal tools to understand the design. Even if the designer builds their understanding of the design purely from documentation, that documentation contains descriptions of the dynamic behavior of the design such as timing diagrams and waveforms.
We present Duet: a general methodology for developing Design Understanding via Experimentation and Testing. Duet that mimics how hardware design experts develop an understanding of complex designs: not just through reading the RTL, but via iterative experimentation. Duet enables AI agents to generate hypotheses about a design, and equips it with tools to test and confirm or refine these hypotheses.
Let’s consider a simple example. Imagine we task an AI agent with describing a design with a reasonably complex finite state machine—for example, an implementation of $\text{I}^{2}\text{C}$. Specifically, let’s consider how an AI agent describes the $\text{I}^{2}\text{C}$ TX path clock stretching feature when it is given access to the RTL, but is not given any experimental tools. Below, we present a snippet of the output produced by such an agent on the left:
Snippet of initial feature description:
Preconditions:
  • target_enable_i = 1
  • rw_bit_q = 1, address
    ACK phase has completed
Snippet of simulation experiment run by agent:
[35] Addr phase done.
[65] Expect sda_o=0: sda_o=1
[95] Expect stretch=1, scl=0:
event_tx_stretch_o=0, scl_o=1
Snippet of note logged by agent:
Clock stretch failed; likely raised SCL too early. Next: keep SCL low longer to allow tHoldData to elapse.
Snippet of fixed feature description:
Preconditions:
  • target_enable_i = 1
  • rw_bit_q = 1, address ACK
    phase has completed
  • keep scl_i low long enough for thd_dat_i during AddrAckWait/Hold before releasing high for AddrAckPulse.
Figure 1: Comparing illustrative snippets of feature-descriptions produced by a Baseline LLM flow (left) and the same flow enhanced with Duet (right), along with snippets of Duet’s experimentation and failure analysis (center).
As we will soon see, this description is missing a key detail. Imagine we use this description as a starting point for a downstream task, such as having an AI agent verify the TX path clock stretching feature. The agent will try and fail to verify the feature based on the description alone.
Now imagine we use Duet to enhance the process and produce deeper design understanding by giving the AI agent access to experimentation tools and prompting it with experimentation strategies when it generates its description. The agent decides to test its initial description using a simulation experiment. It generates a simulation testbench and calls the provided simulation tool, resulting in the log shown in the middle column above. These logs are sent back to the agent. As we can see from the log, the agent encodes its hypotheses in the testbench itself; in this case, the agent’s hypothesis (that its inputs would produce TX clock stretching behavior) proved to be false. The agent analyzes this result, as shown in the second snippet in the middle column. The agent uses the analysis to fix the testbench in later iterations, and successfully exercises the feature (not shown).
This example demonstrates the key insight of this paper: despite the complexity of RTL, AI agents can learn deep design behaviors from simple trial-and-error experimentation.
Duet is a very general tactic that can be applied in any AI workflow operating over code to which experimentation tools can be applied. For example, Duet could be used in documentation generation to produce deeper insights into specific features of the design, or it could be used in formal verification to examine counterexamples to properties which the agent is struggling to prove. Though it is out of the scope of this paper (and this venue), it is likely even useful outside of hardware and EDA applications.
In our evaluation, we show that utilizing Duet leads to better performance on AI-assisted EDA tasks—in the case of our evaluation, formal verification. We compare a baseline flow, which uses an AI agent without access to experimentation tools, to a Duet-enabled flow. Using experimentation, the Duet-enabled flow is better able to understand the design and debug verification failures.
The rest of this paper proceeds as follows. In section II, we describe the implementation of Duet. In section III, we evaluate Duet against a baseline AI workflow the task of formal verification. In section IV, we present related work. In section V we conclude.
II Duet
In this section we provide details on Duet itself. We present one potential implementation of Duet, describe the experimentation tools we have used, and cover pitfalls in implementation.
II-A Duet as a Subroutine
While Duet is an intentionally generic methodology with many possible implementations, in this section we provide the implementation of a “canonical” form of Duet: the DoExperimentation procedure. DoExperimentation implements the heart of Duet: the experimentation loop. Given the design, an experiment description (“expt”), and a set of tools, DoExperimentation carries out the experiment using the provided set of tools:
procedure DoExperimentation(design, expt, tools)
  $\text{tools}\leftarrow\text{tools}+[\textsc{EndExperimentation}]$
  $\text{messages}\leftarrow[\text{``given the following design:''}+\text{design}+\text{``run the following experiment:''}+\text{expt}]$
[…]
   $\text{response}\leftarrow\textsc{LLM}(\text{messages},\text{tools})$; $\text{output}\leftarrow\text{run tool in response}$; $\text{messages}\leftarrow\text{messages}+[\text{response},\text{output}]$
   if called EndExperimentation then break from loop
  end loop
  $\text{report}\leftarrow\textsc{LLM}(\text{messages}+[\text{``generate an experiment report using the previous context''}])$; return report
end procedure
The AI agent has full control over what experiment it runs in each iteration and when to end experimentation. At the end of experimentation, the agent produces an experiment report describing the hypotheses made, the tests run, and the conclusions drawn from the results. Note that modern AI agent programming paradigms (such as the OpenAI Agents SDK [1]) make it easy to implement this experimentation loop; you simply need to provide the prompt and tools, and the loop will be handled internally.
DoExperimentation is a flexible base implementation of Duet in that it takes a text description of the larger experiment to be performed as input and returns a text report of the results as output. Starting from this base, a number of features could be added, e.g. taking other inputs (such as error logs from formal tools or PPA/timing results from backend tools) or producing richer outputs (such as working testbenches demonstrating the discoveries made during experimentation).
II-B Tools
Duet can be flexibly extended to use many tools. Thus far, we have tried simulators, formal verification tools, waveform viewers, and tools for design analysis/signal extraction. In addition, we have also used tools which are themselves agents—see the counterexample replication tool below. What follows is a description of the tools we have used.
Simulation
The simulation tool allows the agent to run a testbench in simulation and see the printed logs. The agent also sees any errors produced by the simulator, in the case of a malformed testbench. While the agent is free to write whatever testbench it would like, our prompts include various tips about testbench structure and common SystemVerilog pitfalls. For example, it is useful to instruct the agent not to use assertions, which may fail and terminate the testbench early, but instead to use copious print statements to produce rich logs.
In our evaluations, we noticed that AI agents seem to rely most heavily on simulation-based experimentation. There are a number of possible explanations. First, even for hardware designers, simulation is straightforward to understand when compared to other tools. Second, there is likely a bias towards simulation in the training data of LLMs, as simulation-based tests are presumably far more common in open-source repositories than formal testbenches.
Waveform viewer
Waveform viewing tools can be a useful supplement to simulation-based experimentation. Waveforms produced by simulation experiments are generally too large to include in the response to the LLM; however, given access to a tool like vcdcat, agents can decide which portions of the waveform to view.
Formal
Formal verification tools such as Jasper give the agent the ability to prove properties or check whether certain states are reachable using coverpoints. As we show in section III, the agent generally does not use formal verification for experimentation, though we suspect that with better prompt engineering and few-shot-examples, the agent could be tuned to utilize formal verification more heavily.
Counterexample replication
The counterexample replication tool is itself an AI agent, which takes a formal property and a counterexample found by a formal verification tool and converts the counterexample into a SystemVerilog simulation testbench which replicates the counterexample. Each call to the counterexample replication tool generally involves 3–6 calls to the simulation tool, in which the agent attempts to produce a working replication. As we see in section III, this proves to be a very valuable tool.
Signal extraction and design analysis
We have also experimented with using tools for signal extraction and design analysis. For example, the Yosys [2] open-source synthesis tool provides a number of useful utilities such as FSM detectors. These types of tools can provide useful high-level information in a deterministic manner. However, similar to the formal tool above, they require careful prompting so the agent knows how and when to use them.
II-C Pitfalls
We have seen a few pitfalls when implementing an experimentation agent. As with all agents, experimentation agents are good at finding ways to “cheat”. For example, if not properly constrained, the counterexample replication tool will simply force internal signals of the design such that the counterexample is trivially replicated, rather than replicating it using the module inputs. Strict prompting fixes this issue. The experimentation agent can also be too lax when gathering evidence for a claim, or in the worst case, can even hallucinate results; simply asking the agent to thoroughly cite experiment logs for each of its claims can mitigate this issue.
III Evaluation
To evaluate Duet, we integrate it into one of our existing AI-assisted workflows. Our AI-assisted workflows tackle various EDA tasks specifically in the hardware verification space; in this evaluation, we will focus primarily on our formal verification flow.
III-A Experimental Setup
In this subsection, we describe our two points of comparison: our baseline AI-assisted agentic formal flow (Baseline) and our Duet-enabled flow (Duet).
Baseline
Our baseline AI-assisted formal verification workflow works as follows. Note that this description is high-level as proprietary details are not included.
  1. 1.
    The AI agent summarizes the design by observing the RTL.
  2. 2.
    Based on the design summary, the AI agent develops a verification plan by enumerating properties to verify.
  3. 3.
    For each property:
    1. i.
      The agent generates a formal testbench for the property.
    2. ii.
      The agent calls a tool which runs the formal testbench through Jasper.
    3. iii.
      Viewing the report from Jasper, the agent returns to step (i) and iterates on the formal testbench until every assertion passes or until it hits its iteration limit.
    4. iv.
      The agent returns the final testbench (passing or non-passing).
  4. 4.
    Verification terminates after the agent has processed each property in the verification plan.
The baseline formal flow outputs a set of testbenches verifying the properties set out in the verification plan.
Duet
Our Duet-enabled workflow adapts the baseline workflow with a few key modifications:
  • The agent’s prompt instructs the agent to use experimentation when debugging formal testbench failure. These instructions include example scenarios of when to use experimentation: for example, to reproduce counterexamples reported by Jasper or to attempt verification on sub-properties.
  • The agent is given access to three experimentation tools (a subset of the tools described in section II-B):
    • Formal tool: runs a formal verification testbench and returns the proof status of asserts, coverpoints, and assumes.
    • Simulation tool: runs a simulation testbench and returns the logs to the agent.
    • Counterexample replication tool: a special, more directed case of the simulation tool which uses simulation to reproduce a Jasper counterexample.
Running the workflows
We run the two workflows on the same design. We use the same design summary and verification plan (steps 1 and 2 of Baseline) for both flows, to ensure we can compare the outputs property-by-property. Each flow is only given a single attempt to prove each property. We use GPT-5 for all LLM queries. We use Jasper as our formal verification tool and Verilator as our simulator.
III-B Results and Analysis
Property Baseline Duet Improved # Duet sim expts # Duet repl expts # Duet formal expts
1 unproven proven 0 7 0
2 unproven unproven, refined $\sim$ 0 5 0
3 proven proven 1 4 0
4 proven proven 3 1 1
5 unproven unproven 1 1 0
6 unproven proven 1 1 0
7 vacuous proven 1 1 0
8 unproven unproven 0 1 0
9 proven proven 1 1 0
10 unproven unproven 3 1 1
TABLE I: Baseline vs. Duet performance on the task of formally verifying properties of a design.
Table I shows the results of using both flows to verify the same 10 properties of the design. We can see that the Duet-enhanced flow doubles the number of proven properties to six, over the three proven in the baseline.
In Table I, we classify a property as having been improved if Duet was able to prove a previously unproven or vacuous assertion in the Baseline. Two of the improved cases (1 and 6) were straightforward instances of the baseline flow failing to prove a property, while the Duet-enhanced flow successfully proved the property. Perhaps more interesting were the the other two cases. Property 7 was proven but vacuous in the baseline. However, via experimentation, the Duet-enabled flow was able to find a non-vacuous refinement of the property and prove it. Property 2 was unproven by the baseline, but the Duet-enabled flow successfully refined it. While not successfully proving the property, experimentation still uncovered a path towards progress (hence the $\sim$).
The last three columns of table I show the number of calls to each of the experimentation tools. Recall that each call to the counterexample replication tool (indicated by the “# repl expts” column) actually represents multiple calls (generally 1-5) to the simulation experimentation tool.
The counterexample replication tool proved to be the most useful when attempting to make progress on formal verification. Every improved property used the counterexample replication tool. In the extreme case of property 1, the agent replicated seven separate counterexamples using the tool. It is perhaps unsurprising that this tool proved the most useful; understanding counterexamples is core to the process of formal verification.
The simulation tool was generally useful as well, though not used on all improved properties. Our general observation is that simulation is favored over formal in a way that suggests the LLM better understands how to craft SystemVerilog simulation testbenches, vs. formal testbenches. The formal tool was rarely used—only in two cases. There are numerous other tools we didn’t include in this evaluation which are treated largely the same by the agent—i.e., mostly ignored. We suspect this could be fixed using better prompting, including clear examples of when each tool might be useful. Additionally, the “formal experimentation tool” is functionally identical to the tool which the agent uses to invoke Jasper on its output testbench, though the name is different and the prompt contains different instructions for using the tool. It may be the case that the agent gets confused by the presence of both tools. Indeed, when the agent iterates on the final testbench by writing a testbench, calling the formal tool, and viewing the logs, it is simply a form of experimentation.
IV Related Work
IV-A LLMs for Software Test Generation
Large Language Models (LLMs) have shown remarkable progress in software test generation. Most existing frameworks for LLM-based software-testing utilize a multi-step pipeline comprised of code-understanding, followed by test-generation, and finally a correction loop where testbenches are fixed based on compiler-feedback. For instance, frameworks including TestSpark [3] for Java and TestPilot [4] for Javascript/Typescript follow such a generation-validation-repair pipeline. ChatUnitTest [5] additionally utilizes an adaptive focal context generation module for modular code-understanding. Similarly, ChatTester [6] also leverages ChatGPT to first understand the focal method, and was able to improve both the compilability and quality of generated assertions compared to direct application of ChatGPT. In the TestART framework [7], the final feedback loop also includes a coverage report – prompting the model to update the testbench to also fill coverage holes. Other methods such as HITS [8] have improved coverage through method slicing, achieving higher branch and line coverage on complex Java benchmarks compared to EvoSuite [9]. An alternative approach to coverage maximization utilized by methods such as MuTAP [10] leverages mutations to maximize coverage by injecting bugs into the code and prompting the LLM to update testbenches to cover any uncaught bugs. There has also been recent work that uses fine-tuning and reinforcement learning to improve LLM’s reasoning abilities for better software testing. For example, [11] fine-tune models to minimize test-smells in unit-tests, and [12] co-evolve the training of code-generation and unit-test-generation to improve performance for both tasks.
Agentic frameworks such as ChatDev [13], and SWE-Agent [14] introduced iterative LLM–tool loops, where models perform reasoning, invoke compilers or linters, and refine outputs based on tool feedback. Recent extensions to verification, like Prompt. Verify. Repeat. [15], adopt similar human-in-the-loop workflows for hardware, coupling LLM reasoning with EDA tools (simulators, formal analyzers) to validate and correct generated artifacts. These “centaur” systems demonstrate that tool-driven iteration can improve reliability but also highlight a lack of structured experimentation frameworks—LLMs must reason not only over text but over the dynamic state of design execution.
The successful application of LLMs in software testing – particularly agentic LLM systems – demonstrates the promise of the same for hardware verification. However, this domain presents some unique challenges that warrant domain-specific solutions. As discussed earlier, RTL code lacks the explicit structural cues that software languages provide, making it difficult for LLMs to understand deeper hardware behaviors and resulting in worse performance on verification tasks. Such deeper design-understanding is possible through Duet’s iterative experimentation.
IV-B LLM Agents for Hardware Verification
Recent work has begun exploring LLMs for hardware verification, particularly for formal property generation and testbench synthesis. AssertLLM [16] converts unstructured hardware specifications (text and waveform diagrams) into SystemVerilog Assertions (SVAs) using a two-stage LLM pipeline. It achieves higher syntactic accuracy than GPT-3.5 or GPT-4 on handcrafted benchmarks but still fails to guarantee semantic correctness, often missing temporal operators or misidentifying relevant signals. Similarly, AssertionBench [17] provides a standardized benchmark for evaluating LLMs on real-world SVA and PSL properties, showing that none of the LLM models evaluated could generate valid assertions an average of no more than 44% accuracy whereas up to 63% generated assertions produced CEX and on average up to 33% of generated assertions were syntactically wrong. A crucial bottleneck is obtaining a precise understanding of the design from RTL, which is made challenging by the limited context-windows of LLMs. Existing work tends to circumvent this by splitting designs into hierarchical modules for independent analysis. [18] proposes incremental proof-generation starting from the basic DUT module and expanding to additional modules. Similarly [19], use LLMs to hierarchically parse the DUT to obtain a formal model and invariants, and employ mutation testing to validate the invariants. Following a similar intuition, Duet also breaks down the task of design-understanding into smaller sub-tasks. But going beyond simply parsing the DUT into smaller modules, Duet allows breaking down the design in terms of functionalities, and builds a deep understanding of each functionality through verified outputs from deep experimentation with tools used by real hardware verification engineers.
In simulation-based verification, AutoBench [20] demonstrated that LLMs can generate self-checking testbenches from RTL module descriptions using a hybrid prompt-template framework. The model achieved a 57% improvement in pass rate over baseline LLM models, but relied heavily on handcrafted scaffolding and lacked temporal reasoning. [21] extract a structured representations of design specification, which are used to prompt an LLM to obtain a set of assertions. These assertions are then refined by iteratively prompting the LLM to fix issues identified by a simulator. LAAG-RV [22] follows a similar protocol, but additionally utilizes a ‘signal-synchronization’ step to refine assertions by prompting an LLM to align the assertions with the design’s RTL code.
In addition to using LLMs for verification, various works have also utilized LLMs to generate RTL code [23, 24, 25, 26, 27], and patch bugs in the design based on precomputed or LLM-generated testbenches [28, 29, 30, 25, 31]. Moreover, there has also been work in not just utilizing existing LLMs for verification, but also to fine-tune open-source models to better equip them with hardware specific domain knowledge. ChipNeMo [32] uses various domain-adaptation strategies including adaptive tokenization, pretraining, retrieval, and so on. [33] utilize version-control data to compile a dataset of open-source hardware design bugs and corresponding fixes.
These works confirm that while LLMs can generate hardware artifacts, domain-specific validation and feedback remain critical. Overall, these systems generally treat the LLM as a one-shot generator, not as an interactive agent capable of experimentation. Moreover, obtaining a deep understanding of the DUT is crucial for reliable verification, but limitations including limited-context windows of LLMs has prevented design-understanding techniques that can scale up with the complexity of designs [19, 18]. Approaches such as VerilogReader [34] use LLMs to enhance DUT code with natural language descriptions to provide greater context of intended functionalities in downstream LLM-tasks. Similarly, RTLExplain [35] generates top-level module summaries via a bottom-up, data-dependency-aware method: extracting a module-dependency graph, producing summaries in reversed topological order, and compiling them into a project-level knowledge base. This achieved 37% accuracy with a RAG agent, compared to 27% with conventional RAG. Once again, the design-understanding obtained through such methods is restricted by their exclusive dependence on static RTL and design specifications, whereas Duet takes advantage of powerful hardware-specific tools that allow AI agents to iteratively verify and refine their design-understanding.
V Conclusion
We present Duet, a general methodology to develop comprehensive design understanding capabilities through iterative experimentation. We demonstrate how experimentation with Duet produces deeper design understanding that has concrete, positive effects on downstream tasks—in our case, formal verification.

### v2

@@
 Specifically, let’s consider how an AI agent describes the $\text{I}^{2}\text{C}$ TX path clock stretching feature when it is given access to the RTL, but is not given any experimental tools.
-Below, we present a snippet of the output produced by such an agent on the left:
+Below, in the left column of fig.
+1, we present a snippet of the output produced by such an agent:
 Snippet of initial feature description:
@@
   • keep scl_i low long enough for thd_dat_i during AddrAckWait/Hold before releasing high for AddrAckPulse.
-Figure 1: Comparing illustrative snippets of feature-descriptions produced by a Baseline LLM flow (left) and the same flow enhanced with Duet (right), along with snippets of Duet’s experimentation and failure analysis (center).
+Figure 1: Comparing illustrative design descriptions produced by a baseline LLM flow (left) and the same flow enhanced with Duet (right), along with snippets of Duet’s experimentation and failure analysis (center).
 As we will soon see, this description is missing a key detail.
 Imagine we use this description as a starting point for a downstream task, such as having an AI agent verify the TX path clock stretching feature.
-The agent will try and fail to verify the feature based on the description alone.
+The agent will try and fail to verify the feature based on the description alone due to a lack of detail.
 Now imagine we use Duet to enhance the process and produce deeper design understanding by giving the AI agent access to experimentation tools and prompting it with experimentation strategies when it generates its description.
@@
 The agent uses the analysis to fix the testbench in later iterations, and successfully exercises the feature (not shown).
+This results in the improved feature description in the right column of fig.
+1
 This example demonstrates the key insight of this paper: despite the complexity of RTL, AI agents can learn deep design behaviors from simple trial-and-error experimentation.
@@
 For example, Duet could be used in documentation generation to produce deeper insights into specific features of the design, or it could be used in formal verification to examine counterexamples to properties which the agent is struggling to prove.
-Though it is out of the scope of this paper (and this venue), it is likely even useful outside of hardware and EDA applications.
+Though it is out of the scope of this paper, it is likely even useful outside of hardware and EDA applications.
 In our evaluation, we show that utilizing Duet leads to better performance on AI-assisted EDA tasks—in the case of our evaluation, formal verification.
@@
 In section II, we describe the implementation of Duet.
-In section III, we evaluate Duet against a baseline AI workflow the task of formal verification.
+In section III, we evaluate Duet against a baseline AI workflow on the task of formal verification.
 In section IV, we present related work.
@@
 The counterexample replication tool is itself an AI agent, which takes a formal property and a counterexample found by a formal verification tool and converts the counterexample into a SystemVerilog simulation testbench which replicates the counterexample.
-Each call to the counterexample replication tool generally involves 3–6 calls to the simulation tool, in which the agent attempts to produce a working replication.
-As we see in section III, this proves to be a very valuable tool.
+Each call to the counterexample replication tool generally involves 1–5 calls to the simulation tool, in which the agent attempts to produce a working replication.
+As we see in section III, this proves to be a very valuable tool as the trial-and-error process of counterexample replication seems to produce deeper design understanding.
 Signal extraction and design analysis
@@
 The experimentation agent can also be too lax when gathering evidence for a claim, or in the worst case, can even hallucinate results; simply asking the agent to thoroughly cite experiment logs for each of its claims can mitigate this issue.
+All of these observations are a reminder that it is all too easy to think of modern LLM-based AI agents as “human-like” intelligence.
+While they are incredibly powerful and useful, they also fail in ways most human experts never would.
 III Evaluation
@@
   • The agent is given access to three experimentation tools (a subset of the tools described in section II-B):
-    • Formal tool: runs a formal verification testbench and returns the proof status of asserts, coverpoints, and assumes.
-    • Simulation tool: runs a simulation testbench and returns the logs to the agent.
-    • Counterexample replication tool: a special, more directed case of the simulation tool which uses simulation to reproduce a Jasper counterexample.
+    • Formal tool: runs a formal verification testbench with Jasper and returns the proof status of asserts, coverpoints, and assumes.
+    • Simulation tool: runs a simulation testbench with Verilator and returns the logs to the agent.
+    • Counterexample replication tool: a special, more directed case of the simulation tool which uses simulation via Verilator to reproduce a Jasper counterexample.
 Running the workflows
-We run the two workflows on the same design.
+We run the two workflows on the same design: a simple round-robin arbiter.
 We use the same design summary and verification plan (steps 1 and 2 of Baseline) for both flows, to ensure we can compare the outputs property-by-property.
+We arbitrarily limit the verification plan to a total of ten properties.
 Each flow is only given a single attempt to prove each property.
-We use GPT-5 for all LLM queries.
+We use GPT-5.0 for all LLM queries.
 We use Jasper as our formal verification tool and Verilator as our simulator.
@@
 Two of the improved cases (1 and 6) were straightforward instances of the baseline flow failing to prove a property, while the Duet-enhanced flow successfully proved the property.
-Perhaps more interesting were the the other two cases.
+Perhaps more interesting were the other two cases.
 Property 7 was proven but vacuous in the baseline.
@@
 Indeed, when the agent iterates on the final testbench by writing a testbench, calling the formal tool, and viewing the logs, it is simply a form of experimentation.
+The evaluation also highlights where Duet can be improved.
+For example, despite careful prompting, there were still some cases where the agent found ways to “cheat”—for example, replicating counterexamples by simply forcing the outputs to the expected values, instead of only manipulating the inputs.
+Interestingly, though, some of these cases still seemed to help the agent with its understanding of the design, suggesting that simply reformatting the counterexamples into a more comprehensible format is itself useful.
+Case Study: Property 1
+To peek inside the Duet flow, we will look at flow artifacts for Property 1.
+Both the baseline and Duet-enabled flows begin by an RTL-only analysis of the design, in which the agent reads the design files and writes out a design summary, which we show in truncated form below:
+Design summary (as implemented):
+- Module and parameters
+  - Top: arbiter
+  - Parameters:
+    - NUM_PORTS: number of request/grant ports ...
+  - Ports:
+    - clk: clock, rising-edge
+    - rst: synchronous reset (active high) ...
+- Core state and combinational signals ...
+- Sequential behavior per clock edge ...
+- What to verify (suggested properties) ...
+As this summary comes only from viewing the RTL, it may be incorrect. It is on the verifier agent to realize this and refine its properties accordingly.
+We then begin the actual process of verification with a prompt:
+Property 001: grant_subset_of_request
+Goal: Prove that in every cycle, grant is a subset of request:
+  (grant & ~request) == 0.
+As we will soon see, this property isn’t correct as stated; however, the agent will be able to discover this and refine the property via experimentation.
+After some reasoning, the agent produces its first attempt at a formal testbench.
+We see this behavior often—even when experimentation tools are available, the agent will often attempt to achieve the goal without them first.
+This behavior can be controlled via prompting, but we have found it more beneficial to focus on ensuring that the agent uses the experimentation tools if and when its initial attempts fail.
+The following is a truncated version of its initial attempt at a formal testbench:
+module arbiter_grant_subset_sva #(...) (...);
+  property p_grant_subset;
+    @(posedge clk) (grant & ~request) == ’0;
+  endproperty
+  assert property (p_grant_subset);
+endmodule
+When the agent calls Jasper on this initial attempt at a formal testbench, Jasper produces counterexamples.
+The agent then uses the counterexample replication tool.
+Recall that the counterexample replication tool is itself an agent which uses multiple calls to the simulation tool to replicate a given counterexample.
+Below we show the truncated output of the counterexample replication tool:
+module tb_arbiter_ce;
+  arbiter #(...) dut (...);
+  initial clk = 1’b0; always #5 clk = ~clk;
+  int cyc; initial cyc = 0; always @(posedge clk) cyc <= cyc + 1;
+  // Basic stimulus to steer token to index 4 and produce a grant on bit 4
+  initial begin ... end
+  // Create and check the counterexample right when grant[4] goes high
+  always @(posedge clk) begin
+    if (cyc == 2) begin
+      rst     <= 1’b1; request <= 6’h22; #1;
+      assert (grant[4] == 1’b1 && request[4] == 0);
+    end
+  end
+endmodule
+Though the agent doesn’t directly use the output of the counterexample replication tool, it seems that the process of replication gives the agent a deeper understanding of the design.
+In this case, it shows the agent that the timing behavior of the given property is more complex than originally proposed, and helps the agent refine the property.
+Finally, the agent is able to produce a formal testbench which passes:
+module arbiter_grant_subset_sva #(...) (...);
+  property p_grant_subset_prev_req;
+    @(posedge clk)
+      1’b1 |-> ##1 ((grant & ~\$past(request)) == ’0);
+  endproperty
+  assert property (p_grant_subset_prev_req);
+endmodule
+Compared to the Duet-enabled flow, the baseline flow is unable to discover the exact timing of the property.
+While it iterates a number of times on the formal testbench, it fails to pull deeper meaning from the counterexamples returned by Jasper.
 IV Related Work
@@
 The model achieved a 57% improvement in pass rate over baseline LLM models, but relied heavily on handcrafted scaffolding and lacked temporal reasoning.
-[21] extract a structured representations of design specification, which are used to prompt an LLM to obtain a set of assertions.
+[14] extract a structured representation of design specifications, which are used to prompt an LLM to obtain a set of assertions.
 These assertions are then refined by iteratively prompting the LLM to fix issues identified by a simulator.
@@
 Overall, these systems generally treat the LLM as a one-shot generator, not as an interactive agent capable of experimentation.
-Moreover, obtaining a deep understanding of the DUT is crucial for reliable verification, but limitations including limited-context windows of LLMs has prevented design-understanding techniques that can scale up with the complexity of designs [19, 18].
+Moreover, obtaining a deep understanding of the DUT is crucial for reliable verification, but limitations including limited-context windows of LLMs have prevented design-understanding techniques that can scale up with the complexity of designs [8, 18].
 Approaches such as VerilogReader [12] use LLMs to enhance DUT code with natural language descriptions to provide greater context of intended functionalities in downstream LLM-tasks.
@@
 We present Duet, a general methodology to develop comprehensive design understanding capabilities through iterative experimentation.
+We present a generally-applicable implementation of Duet which is useful in many potential AI-assisted flows.
 We demonstrate how experimentation with Duet produces deeper design understanding that has concrete, positive effects on downstream tasks—in our case, formal verification.